# What Are They Doing? Joint Audio-Speech Co-Reasoning

### v1

@@ Abstract
In this paper, we investigate how well ALLMs can perform joint audio-speech processing. Specifically, we introduce Joint Audio-Speech Co-Reasoning (JASCO), a novel task that unifies audio and speech processing, strictly requiring co-reasoning across both modalities. We release a scene-reasoning dataset called ”What Are They Doing” and establish a joint audio-speech benchmark to evaluate the joint reasoning capability of popular ALLMs. Additionally, we provide deeper insights into the models’ behaviors by analyzing their dependence on each modality.
@@ I Introduction
Compared to the unification of audio and speech accomplished at the model architecture level, from the task perspective, almost all the tasks used for training and evaluating ALLMs are still limited to a single modality. This includes traditional audio/speech tasks as well as open-ended question-answering tasks such as Audio/Speech Question Answering (AQA/SQA). From the training perspective, single-modality tasks cannot allow for the development of a truly intelligent audio-speech LLM that can achieve joint reasoning, and they may also lead to known problems such as task overfitting [14]. From the evaluation perspective, the established single-modality benchmarks [23, 24, 25] cannot effectively reflect the model’s joint understanding capacity and cannot determine if the model’s reasoning is misled by its preference for a certain modality.
@@
  • There is dominant information in one certain modality that is sufficient for making good reasoning, even without the need for the other modality (In Sample 2, the answer can be directly inferred from the audio even without using speech).
@@ II-A Task Design
We established several constraints to guarantee that the proposed task qualifies as a real joint task and strictly performs co-reasoning:
@@ III-B Audio Preparation
Due to the complexity of the task and the significant manual effort required, we were able to produce 80 high-quality co-reasoning samples. Although this quantity is insufficient for model training, it can serve as a unique test set for ALLM evaluation to demonstrate their potential abilities and behaviors.
@@
The average model-as-judge scores calculated for four popular ALLMs based on their best responses to each sample.
model Best-Mean (between 0 and 2)
WavLLM-7B 0.49
LTU-AS-7B 1.04
SALMONN-7B 1.21
Qwen2-Audio-Instruct-7B 1.38
@@ IV-A Experimental Setup
We benchmarked four popular ALLMs on the proposed What Are They Doing dataset: WavLLM-7B [^3] https://github.com/microsoft/SpeechT5/tree/main/WavLLM, LTU-AS-7B [^4] https://github.com/YuanGongND/ltu/tree/main/src/ltu_as, SALMONN-7B [^5] https://huggingface.co/tsinghua-ee/SALMONN-7B/tree/main, Qwen2-Audio-Instruct-7B [^6] https://huggingface.co/Qwen/Qwen2-Audio-7B-Instruct [34]. It should be mentioned that almost all the training data of WavLLM are speech-related, so it can be considered a pure speech LLM and is used as a comparison to other models. For each sample, we tested with 8 different instruction prompts to reduce the bias of the evaluation [35]. In our experiments, we found that the models did not always make a prediction on the speaker’s action. To reduce the difficulty for the models, we included step-by-step thinking [36] instructions by asking them to first detect the audio sounds and transcribe the spoken texts. An example can be: ”What sound can you hear? What does the speaker say? Guess what activity the speakers are engaged in using both the audio sound and the spoken text.”.
We calculated the results using the Model-As-Judge method mentioned in section 2.C. Considering both overall performance and accessibility, we selected Llama-3.1-70B-Instruct  [^7] https://huggingface.co/meta-llama/Meta-Llama-3.1-70B-Instruct as the model judge, which is a powerful, general-purpose, and open-source LLM.
@@ IV-B Results and Analysis
In Table 2, we present the average scores of each model on the proposed dataset. In order to reflect the upper limit of the models’ capabilities, for each sample in the dataset, we selected the best score among all the 8 predictions for computing the average, which is referred to as Best-Mean. As can be seen from the results, Qwen2 achieved the highest score of 1.38, followed by SALMONN at 1.21 and LTU-AS at 1.04. In contrast, WavLLM only obtained a score of 0.49, suggesting that using only speech leads to poor performance on the task, which also confirms the validity of the proposed dataset. Besides, all the models exhibit a substantial gap below the highest score 2, indicating their limited co-reasoning power under the proposed task.
Figure 2: Modality-dependence results of 4 popular ALLMs. Red, blue, and green are used to represent the proportion of audio-dependent (A), speech-dependent (S), and both-dependent (A-S) responses.
In Figure 2, we present the results of the models’ modality-dependence. This calculation is based on all the responses obtained using all instruction prompts across all samples. Responses in which the models did not provide predictions regarding the speaker’s actions were excluded from this analysis. As the results indicate, the Qwen2 model still leads with a both-dependent percentage of 56%, followed by LTU-AS at 49% and Salmonn at 48%. In contrast, WavLLM records only 22% but has the highest speech dependence at 74%, which highlights its nature as a speech LLM. These results indicate that the current ALLMs lack a consistent practice of joint reasoning. We further discover that despite the equal importance of audio and speech in the JASCO task, both Qwen2 and LTU-AS lean more toward being speech-LLMs than audio-LLMs with always higher speech-dependent percentage than audio-dependent percentage. Unlike the others, Salmonn demonstrates a more balanced dependence on audio and speech. Through our benchmark, we hope to encourage more research on joint audio-speech tasks and well-designed joint audio-speech datasets.
@@ V Conclusion
In this work, we propose the Joint Audio-Speech Co-Reasoning task which focuses on performing joint reasoning based on both audio and speech modalities. Additionally, we introduce the What Are They Doing dataset, a freely available open-ended scene-reasoning style dataset. We build a joint audio-speech benchmark for the existing popular ALLMs and reveal their modality-dependence. Through our work, we hope to provide deeper insights into the behavior of existing ALLMs and lay the groundwork for the future development of truly intelligent audio-speech LLMs.

### v2

@@ Abstract
In this paper, we establish a novel benchmark to investigate how well ALLMs can perform joint audio-speech processing. Specifically, we introduce Joint Audio-Speech Co-Reasoning (JASCO), a novel task that unifies audio and speech processing, strictly requiring co-reasoning across both modalities. We also release a scene-reasoning dataset called ”What Are They Doing”. Additionally, we provide deeper insights into the models’ behaviors by analyzing their dependence on each modality.
@@ I Introduction
Compared to the unification of audio and speech accomplished at the model architecture level, from the task perspective, almost all the tasks used for training and evaluating ALLMs are still limited to a single modality. This includes traditional audio/speech tasks as well as open-ended question-answering tasks such as Audio/Speech Question Answering (AQA/SQA). From the training perspective, single-modality tasks cannot allow for the development of a truly intelligent audio-speech LLM that can achieve joint reasoning, and they may also lead to known problems such as task overfitting [14]. From the evaluation perspective, the single-modality benchmarks [23, 24, 25] cannot effectively reflect the model’s joint understanding capacity and cannot determine if the model’s reasoning is misled by its preference for a certain modality.
@@
  • There is dominant information in one certain modality that is sufficient for good reasoning, even without using the other modality (In Sample 2, the answer can be directly inferred from the audio even without using speech).
@@ II-A Task Design
We established several constraints to guarantee that the proposed task strictly performs co-reasoning:
@@ III-B Audio Preparation
Due to the complexity of the task and the significant manual effort required, we were able to produce 80 high-quality co-reasoning samples. While this quantity is insufficient for model training, it can serve as a unique test set for ALLM evaluation to demonstrate their potential abilities and behaviors.
@@
[…]
LLM-Judge ALLM WavLLM LTU-AS SALMONN Qwen2-Audio
Llama-3.1-70B-Instruct 0.49 1.04 1.21 1.38
Qwen2.5-72B-Instruct 0.33 1.19 1.29 1.43
Mistral-123B-Instruct 0.13 0.74 0.80 0.88
Average 0.31 0.99 1.10 1.23
@@ IV-A Experimental Setup
We benchmarked four popular ALLMs on the proposed What Are They Doing dataset: WavLLM-7B [^3] https://github.com/microsoft/SpeechT5/tree/main/WavLLM, LTU-AS-7B [^4] https://github.com/YuanGongND/ltu/tree/main/src/ltu_as, SALMONN-7B [^5] https://huggingface.co/tsinghua-ee/SALMONN-7B/tree/main, Qwen2-Audio-Instruct-7B [^6] https://huggingface.co/Qwen/Qwen2-Audio-7B-Instruct [34]. It should be mentioned that almost all the training data of WavLLM are speech-related, so it can be considered a pure speech-LLM and is used as a comparison to other models. For each sample, we tested with 8 different instruction prompts to reduce the bias of the evaluation [35]. In our experiments, we found that the models did not always make a prediction on the speaker’s action. To reduce the difficulty for the models, we included step-by-step thinking [36] instructions by asking them to first detect the audio sounds and transcribe the spoken texts. An example can be: ”What sound can you hear? What does the speaker say? Guess what activity the speakers are engaged in using both the audio sound and the spoken text.”.
We scored the ALLMs’ predictions using the Model-As-Judge method mentioned in section 2.C. In order to enhance the evaluation’s reliability, we selected three open-sourced LLMs as judges and calculated their average. The selected judges are Llama-3.1-70B-Instruct  [^7] https://huggingface.co/meta-llama/Meta-Llama-3.1-70B-Instruct, Qwen2.5-72B-Instruct [^8] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct and Mistral-123B-Instruct [^9] https://huggingface.co/mistralai/Mistral-Large-Instruct-2411.
@@ IV-B Results and Analysis
In order to reflect the upper limit of the ALLMs’ capabilities, for each sample, we selected only the best score among the 8 predictions for computing the average across the entire dataset, which is referred to as Best-Mean. For each ALLM, we ran three LLM judges and calculated the average Best-Mean among the judges. As can be seen from Table 2, Qwen2-Audio achieved the highest average Best-Mean of 1.23, followed by SALMONN at 1.10 and LTU-AS at 0.99. In contrast, WavLLM only obtained a score of 0.31, suggesting that training with only speech leads to poor performance on the proposed benchmark, which also confirms the validity of the proposed dataset. Besides, all the models exhibit a substantial gap below the highest score 2, indicating their limited co-reasoning power under the proposed JASCO task.
Figure 2: Modality-dependence results of 4 popular ALLMs. The four sub-figures represent evaluations from three LLM judges and their average. Red, blue, and green are used to represent the proportion of audio-dependent (A), speech-dependent (S), and both-dependent (A-S) responses.
In Figure 2, we present the results of the ALLMs’ modality-dependence. The calculation is based on all the responses obtained using all 8 instruction prompts across all samples. Responses in which the ALLMs did not provide predictions regarding the speaker’s actions were excluded from this analysis. Similarly, we report the results from the three judges as well as their average. As the results indicate, the Qwen2-Audio still leads with an average both-dependent percentage of 50%, followed by Salmonn at 42% and LTU-AS at 40%. In contrast, WavLLM records only 11% but has the highest average speech-dependence at 77%, which highlights its nature as a speech LLM. It can be seen from the results that the current ALLMs still lack a consistent practice of joint reasoning. We further discover that despite the equal importance of audio and speech in the JASCO task, both Qwen2 and LTU-AS lean more toward being speech-LLMs than audio-LLMs with always higher speech-dependent percentage than audio-dependent percentage. Unlike the others, Salmonn demonstrates a more balanced dependence on audio and speech. By presenting this benchmark, we aim to provide diagnostic insights into the behavior of existing ALLMs.
@@ V Conclusion
In this work, we built a joint audio-speech benchmark for ALLMs and revealed their modality-dependence. Specifically, we propose the Joint Audio-Speech Co-Reasoning task and the What Are They Doing dataset. Through our work, we hope to provide deeper insights into the behavior of existing ALLMs and lay the groundwork for the future development of truly intelligent audio-speech LLMs.